# Minimal Construction of Graphs with Maximum Robustness

### v1

@@ 3.2 Case 2: Graphs with Even $n$
We can continue this process of (i) drawing edges between $v_{i}\in\mathcal{S}_{1}$ and $\gamma$ nodes in $\mathcal{S}_{2}$ and (ii) moving $v_{i}\in\mathcal{S}_{1}$ into $\mathcal{S}_{2}$ until $\mathcal{S}_{1}$ becomes empty. Then, we can do these steps $\gamma-1$ times, as we initially have $|\mathcal{S}_{1}|=\gamma-1$. In other words, for each $\gamma-1$ pair of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, we require $\gamma$ additional edges. Since we have started from the setting where $\mathcal{S}_{2}=\mathcal{V}_{S}$, we have $|\mathcal{E}|\geq|\mathcal{E}_{S}|+\gamma(\gamma-1)$. Therefore, if $\gamma$ is even, $|\mathcal{E}|\geq\frac{\gamma^{2}+2}{2}+\gamma(\gamma-1)=\frac{\gamma(3\gamma-2)+2}{2}$. If $\gamma$ is odd, and $|\mathcal{E}|\geq\frac{\gamma^{2}+1}{2}+\gamma(\gamma-1)=\frac{\gamma(3\gamma-2)+1}{2}$.
Remark 3.12.
While the primary goal of this section has been to establish bounds on the edge counts for $\gamma$-robust graphs, we also have identified some interesting graph structures that are necessary for $\gamma$-robust graphs. These structures arise solely from the definition of $r$-robustness (Definition 6) and provide topological representations of what $r$-robustness requires. A more systematic exploration of these necessary structures could yield valuable insights, particularly into how to relax the often dense connectivity requirements associated with robustness. Investigating this direction further presents a promising avenue for future work.
From Lemma 3.10, we can obtain a corollary similar to Corollary 3.4:
Corollary 3.13.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an $r$-robust graph, where $r\in\mathbb{Z}_{>0}$ and $|\mathcal{V}|=n$ is even. Then,
$|\mathcal{E}|\geq\left\lfloor{\frac{r(3r-2)+2}{2}}\right\rfloor.$ (5)
Proof 3.14.
In general, the more nodes a graph has, the more edges it must have to maintain a certain robustness. From [30], it can be seen shown that $2r$ is the smallest even number of nodes a graph to have to be $r$-robust. In other words, $r$-robustness is the maximum robustness a graph with $2r$ nodes can achieve. Then by Lemma 3.10, a graph of even number of nodes needs at least $\left\lfloor{\frac{r(3r-2)+2}{2}}\right\rfloor$ edges to be $r$-robust.
Remark 3.15.
We compare the bounds established in Lemmas 3.2 and 3.10 to observe that the lower bound in the latter is strictly stronger. In other words, while the latter also applies to $\gamma$-robust graphs with even $n$, the former offers a tighter condition. This is also reflected in Corollary 3.4 and 3.13, where the latter provides a stronger necessary condition on the edge counts for any graphs with even $n$ to become $r$-robust than the former does.
@@ 4 Construction of $\gamma$-MERGs
In the previous section, we have analyzed and established the lower bounds on edge counts that graphs of $n$ nodes must satisfy in order to achieve $\gamma$-robustness (where $\gamma=\lceil\frac{n}{2}\rceil$ is the maximum robustness). Proceeding with the lower bounds, we construct $\gamma$-MERGs, or graphs that provably achieve $\gamma$-robustness with a minimal set of edges. This is formally given in Theorem 4.16, as our first main result:
Theorem 4.16.
@@
Proof 4.17.
To prove that $\mathcal{G}$ is $\gamma$-MERG, we need to prove two things for both odd and even values of $n$: (i) it is $\gamma$-robust and (ii) its edge count $|\mathcal{E}|$ equals to the lower bounds specified in Lemma 3.2 or Lemma 3.10, depending on the value of $n$.
@@
2) If $\mathcal{X}\cap\mathcal{S}_{1}\neq\emptyset$, denote $|\mathcal{X}\cap\mathcal{S}_{1}|=p\leq\gamma-1$. Then, we have $1\leq|\mathcal{X}^{\prime}\cap\mathcal{S}_{1}|\leq\gamma-1-p$, $|\mathcal{X}\setminus\mathcal{S}_{1}|=\gamma+1-p$, and $p-1\leq|\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}|\leq\gamma-2$. For $i\in\mathcal{X}\cap\mathcal{S}_{1}$, $|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}|=|\mathcal{X}\setminus\{i\}|-(p-1)=\gamma-(p-1)$, as $\mathcal{X}$ forms a $(\gamma+1)$-clique. For $p=1$, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq|\mathcal{X}\setminus\{i\}|=\gamma$. For $p\geq 2$, note that by construction, every node in $\mathcal{X}^{\prime}$ has edges with any $\gamma$ nodes in $\mathcal{X}$, where $|\mathcal{X}|=\gamma+1$. Speaking differently, every node in $\mathcal{X}^{\prime}$ does not have an edge with only one node in $\mathcal{X}$. Thus, all nodes in $\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}$ have edges with at least $p-1$ nodes in $\mathcal{X}\cap\mathcal{S}_{1}$, because it may not have an edge with only one node in $\mathcal{X}\cap\mathcal{S}_{1}$. Then, as $p-1\leq|\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}|\leq\gamma-2$, there are at least a total of $(p-1)(p-1)$ edges between the nodes in $\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}$ and $\mathcal{X}\cap\mathcal{S}_{1}$. Since $|\mathcal{X}\cap\mathcal{S}_{1}|=p$ and by Pigeonhole Principle, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}^{\prime}|\geq\big{\lceil}{\frac{(p-1)(p-1)}{p}}\big{\rceil}=\big{\lceil}{\frac{p^{2}-2p+1}{p}}\big{\rceil}=p-1$. Then, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|=|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}|+|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}^{\prime}|\geq\gamma-(p-1)+p-1=\gamma$. Thus, whether $\mathcal{S}_{1}$ contains a node in $\mathcal{X}$ or not, $\mathcal{S}_{1}$ is always $\gamma$-reachable.
Now we show that $\mathcal{E}$ is a minimal set. $\mathcal{G}$ has $\frac{(\gamma+1)(\gamma)}{2}$ edges to form a $(\gamma+1)$-clique. Additionally, we introduce $\gamma(\gamma-2)$ edges for connecting each of the $\gamma-2$ nodes in $\mathcal{X}^{\prime}$ with $\gamma$ nodes in $\mathcal{X}$. Summing them up, we have $\frac{3\gamma(\gamma-1)}{2}$ which equals the lower bound given in Lemma 3.2.
@@
2) If $\mathcal{X}\cap\mathcal{S}_{1}\neq\emptyset$, let a node $i\in\mathcal{X}\cap\mathcal{S}_{1}$. Note that after removing an edge among $\lceil\frac{\gamma-2}{2}\rceil$ pairs of nodes in $\mathcal{X}$, there exists at least one (two in case of even $\gamma$) node(s) that is connected to all other nodes in $\mathcal{G}$. In other words, $\exists v_{*}\in\mathcal{V}$ s.t. $|N_{v_{*}}|=2\gamma-1$. Now, when $11\leq|\mathcal{S}_{1}|\leq\gamma-$ 1, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq\gamma$, making $\mathcal{S}_{1}$ $\gamma$-reachable. In case $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$, we know there exists $v_{*}$ in either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ such that $|\mathcal{N}_{v_{*}}\setminus\mathcal{S}_{i}|\geq\gamma$, making $\mathcal{S}_{i}$ $\gamma$-reachable. Thus, whether $\mathcal{S}_{1}$ contains a node in $\mathcal{X}$ or not, either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ is always $\gamma$-reachable.
Now, we examine its edge set. The $\gamma$ nodes in $\mathcal{X}$ connecting to every other node adds up to a total of $\frac{\gamma(\gamma-1)+2\gamma(\gamma)}{2}=\frac{3\gamma^{2}-\gamma}{2}$ edges. Then, if we subtract $\lceil\frac{\gamma-2}{2}\rceil$ from it, we get $\frac{\gamma(3\gamma-2)+2}{2}$ for even $\gamma$ and $\frac{\gamma(3\gamma-\gamma)+1}{2}$ for odd $\gamma$, which equals the lower bound given in Lemma 3.10.
@@
Theorem 4.16 provides a systematic procedure for constructing $\gamma$-MERGs, i.e., graphs that achieve maximum $r$-robustness with a minimal set of edges, giving a solution to Problem 1. As explicitly stated in the theorem, $\mathcal{G}$ contains an $(r+1)$-clique when $n$ is odd, further demonstrating Lemma 1. For even $n$, although not explicitly stated, we can observe that $\mathcal{G}$ contains a $\lfloor\frac{\gamma+4}{2}\rfloor$-clique, as given by Lemma 3.10. Furthermore, the theorem suggests that the bounds given in Lemma 3.2 and 3.10 are tight. In fact, one removal of any edge from a $\gamma$-MERG immediately drops the graph’s robustness level by at least one (as demonstrated in Sec. 7.2).
Some examples of $\gamma$-MERGs with $n=9$ (left) and $n=10$ (right) are visualized in Figure 2. The $5$-MERG with $n=9$ actually has a maximum clique of size $6$ (i.e., $\{0,1,2,3,4,5\}$), while the one with $n=10$ contains a $4$-clique as the maximum clique (i.e., $\{0,2,4,7\}$). These examples align with the necessary maximum clique sizes analyzed in Lemma 1 and 3.10: $(\gamma+1)$-cliques for odd $n$ and $\lfloor\frac{\gamma+4}{2}\rfloor$-cliques for even $n$.
@@ 5 Bounds on Edge Counts for $(\gamma,\gamma)$-MERGs
In Section 3 and 4, we have examined and constructed $\gamma$-MERGs to address Problem 1. In the following two sections, we focus on $(\gamma,\gamma)$-MERGs, graphs with a given number of nodes that achieve maximum $(r,s)$-robustness with a minimal set of edges, thereby addressing Problem 2.
Similar to what we have done in Section 3, since $\gamma=\lceil\frac{n}{2}\rceil$, we have two cases: $n=2\gamma-1$ for odd $n$ and $n=2\gamma$ for even $n$. However, $(\gamma,\gamma)$-MERGs for odd $n$ have already been characterized in [4, Lemma 4], which is given below for completeness:
Lemma 5.18.
@@
Lemma 5.18 essentially confirms that when $n$ is odd, the graph must be a complete graph to be a $(\gamma,\gamma)$-MERG. This also implies that when $n$ is odd, for a graph to be $(\gamma,\gamma)$-robust, it must have at least $\frac{(2\gamma-1)(2\gamma-2)}{2}=(\gamma-1)(2\gamma-1)$ edges.
@@ 5.1 Graphs with Even $n$
We now introduce the lower bounds on edge counts for $(\gamma,\gamma)$-robust graphs when $n$ is even. This is formally established in Lemma 5.22. This lemma will be later used in Section 6 to construct $(\gamma,\gamma)$-MERGs for any $n$. However, before that, we introduce another useful lemma from [4]:
Lemma 5.19.
@@
Proof 5.21.
As the direct application of Lemma 6, we can conclude $\delta_{\min}(\mathcal{G})\geq 2(r-1)$. Then, each node has at least $2(r-1)$ neighbors. Thus, since $\mathcal{G}$ is an undirected graph, $|\mathcal{E}|\geq\frac{2n(r-1)}{2}=n(r-1)$.
Corollary 5.20 provides both the minimum degree required for each node and the total number of edges necessary for a graph to be $(r,r)$-robust. Similar to how we have used the notion of cliques as the building blocks of our analyses in Section 3, we will use Corollary 5.20 to analyze graphs with even $n$. By Corollary 5.20, we know for a graph with even $n$ to be $(\gamma,\gamma)$-MERG, it has to have at least $n(\gamma-1)=2\gamma(\gamma-1)$ edges. However, by leveraging the definition of $(r,s)$-robustness (Definition 8), we can derive a tighter lower bound:
Figure 3: This figure visualizes the contradiction shown in the proof for Lemma 5.22. The yellow solid lines represent edges, whereas the colored dotted lines represent the absences of edges. There are two disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$, in which all nodes only have $\gamma-1$ neighbors outside of the subsets, necessitating at least $\lceil\frac{\gamma}{2}\rceil$ of the colored dotted lines to become actual edges to resolve the contradiction.
Lemma 5.22.
@@
Proof 5.23.
By Corollary 5.20, we know $\delta_{\min}(\mathcal{G})\geq 2(\gamma-1)$ and $|\mathcal{E}|\geq(2\gamma)(\gamma-1)$. In the rest of the proof, we will show that some nodes need more than $2(\gamma-1)$ neighbors for $\mathcal{G}$ to satisfy definition of $(r,s)$-robustness (Definition 8). Assume to the contrary that every node $i\in\mathcal{V}$ only has $2(\gamma-1)$ neighbors and $\mathcal{G}$ is $(\gamma,\gamma)$-robust. Because (i) there are $2\gamma$ nodes and (ii) each node has $2(\gamma-1)$ neighbors, we can form disjoint $\gamma$ pairs of nodes that do not have an edge together. Now, form a disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that each subset contains one node of each pair that do not have an edge. Since we have $\gamma$ pairs of nodes without an edge, $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$. By the definition of $(r,s)$-robustness, we need at least $\gamma$ nodes that have $\gamma$ or more neighbors outside of their subsets. However, we have reached a contradiction, as all nodes in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ have only $\gamma-1$ neighbors outside of their subsets. To resolve the contradiction, we need at least $\lceil\frac{\gamma}{2}\rceil$ edges, which are used to create edges between at least $\gamma$ nodes (as shown in Fig 3). Thus, combining everything, we need at least $2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$ edges.
Lemma 5.22 provides the necessary conditions on the edge counts of $(\gamma,\gamma)$-robust graphs with even numbers of nodes. We can also derive an interesting necessary graph structure by examining Lemma 5.22, which we provide below:
Corollary 5.24.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes where $n$ is even, and let $\gamma=\frac{n}{2}$. Then, if $\mathcal{G}$ is $(\gamma,\gamma)$-robust, it must contain a $T(\gamma)$-clique, where
$T(\gamma)<\frac{2\gamma^{2}}{2\gamma-\big{\lceil}{\frac{\gamma}{2}}\big{\rceil}}+1.$ (8)
Proof 5.25.
We begin by recalling Turán’s theorem, which states that undirected graphs of $n$ nodes without a $k$-clique must have at most $(1-\frac{1}{k-1})\frac{n^{2}}{2}$ edges [43]. In other words, if a graph has more than $(1-\frac{1}{k-1})\frac{n^{2}}{2}$ edges, it must contain a $k$-clique.
Then, combining this with Lemma 5.22, $(\gamma,\gamma)$-robust graphs must contain a clique size of $T(\gamma)$, where
$2\gamma(\gamma-1)+\Big{\lceil}{\frac{\gamma}{2}}\Big{\rceil}$ $>\Big{(}1-\frac{1}{T(\gamma)-1}\Big{)}\frac{n^{2}}{2}$
$=2\gamma^{2}\Big{(}1-\frac{1}{T(\gamma)-1}\Big{)}.$
Rewriting the above, we get
$12\gamma-\Big{\lceil}{\frac{\gamma}{2}}\Big{\rceil}<\frac{2\gamma^{2}}{T(\gamma)-1}\to T(\gamma)<\frac{2\gamma^{2}}{2\gamma-\big{\lceil}{\frac{\gamma}{2}}\big{\rceil}}+$ 1 .
Corollary 5.24 characterizes the necessary maximum clique size that any $(\gamma,\gamma)$-robust graphs with even $n$ must contain. Roughly speaking, this implies the necessary maximum clique has to be at least size of $\lfloor{\frac{4\gamma}{3}}\rfloor+1$. Now, we will construct $(\gamma,\gamma)$-MERGs in the next section.

### v2

@@ 3.2 Case 2: Graphs with Even $n$
We continue this process of (i) drawing edges between $v_{i}\in\mathcal{S}_{1}$ and $\gamma$ nodes in $\mathcal{S}_{2}$ and (ii) moving $v_{i}\in\mathcal{S}_{1}$ into $\mathcal{S}_{2}$ until $\mathcal{S}_{1}$ becomes empty. Then, we can do these steps $\gamma-1$ times, as we initially have $|\mathcal{S}_{1}|=\gamma-1$. In other words, for each $\gamma-1$ pair of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, we require $\gamma$ additional edges. Since we have started from the setting where $\mathcal{S}_{2}=\mathcal{V}_{S}$, we have $|\mathcal{E}|\geq|\mathcal{E}_{S}|+\gamma(\gamma-1)$. Therefore, if $\gamma$ is even, $|\mathcal{E}|\geq\frac{\gamma^{2}+2}{2}+\gamma(\gamma-1)=\frac{\gamma(3\gamma-2)+2}{2}$. If $\gamma$ is odd, and $|\mathcal{E}|\geq\frac{\gamma^{2}+1}{2}+\gamma(\gamma-1)=\frac{\gamma(3\gamma-2)+1}{2}$. ∎
[…]
@@ 4 Construction of $\gamma$-MERGs
In the previous section, we have established the lower bounds on edge counts that graphs of $n$ nodes must satisfy in order to achieve $\gamma$-robustness (where $\gamma=\lceil\frac{n}{2}\rceil$ is the maximum robustness). Proceeding with the lower bounds, we construct $\gamma$-MERGs, or graphs that provably achieve $\gamma$-robustness with a minimal set of edges. This is formally given in Theorem 1, as our first main result:
Theorem 1.
@@
Proof.
To prove that $\mathcal{G}$ is $\gamma$-MERG, we need to prove two things for both odd and even values of $n$: (i) it is $\gamma$-robust and (ii) its edge count $|\mathcal{E}|$ equals to the lower bounds specified in Lemma 2 or Lemma 5, depending on the value of $n$.
@@
2) If $\mathcal{X}\cap\mathcal{S}_{1}\neq\emptyset$, denote $|\mathcal{X}\cap\mathcal{S}_{1}|=p\leq\gamma-1$. Then, we have $1\leq|\mathcal{X}^{\prime}\cap\mathcal{S}_{1}|\leq\gamma-1-p$, $|\mathcal{X}\setminus\mathcal{S}_{1}|=\gamma+1-p$, and $p-1\leq|\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}|\leq\gamma-2$. For $i\in\mathcal{X}\cap\mathcal{S}_{1}$, $|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}|=|\mathcal{X}\setminus\{i\}|-(p-1)=\gamma-(p-1)$, as $\mathcal{X}$ forms a $(\gamma+1)$-clique. For $p=1$, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq|\mathcal{X}\setminus\{i\}|=\gamma$. For $p\geq 2$, note that by construction, every node in $\mathcal{X}^{\prime}$ has edges with any $\gamma$ nodes in $\mathcal{X}$, where $|\mathcal{X}|=\gamma+1$. Speaking differently, every node in $\mathcal{X}^{\prime}$ does not have an edge with only one node in $\mathcal{X}$. Thus, all nodes in $\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}$ have edges with at least p1p-1 nodes in $\mathcal{X}\cap\mathcal{S}_{1}$, because it cannot have an edge with only one node in $\mathcal{X}\cap\mathcal{S}_{1}$. Then, as $p-1\leq|\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}|\leq\gamma-2$, there are at least a total of (p1)(p1)$(p-1)$ edges between the nodes in $\mathcal{X}^{\prime}\setminus\mathcal{S}_{1}$ and $\mathcal{X}\cap\mathcal{S}_{1}$. Since $|\mathcal{X}\cap\mathcal{S}_{1}|=p$ and by Pigeonhole Principle, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}^{\prime}|\geq\big\lceil{\frac{(p-1)(p-1)}{p}}\big\rceil=\big\lceil{\frac{p^{2}-2p+1}{p}}\big\rceil=p-1$. Then, $\exists i\in\mathcal{X}\cap\mathcal{S}_{1}$ such that $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|=|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}|+|(\mathcal{N}_{i}\setminus\mathcal{S}_{1})\cap\mathcal{X}^{\prime}|\geq\gamma-(p-1)+p-1=\gamma$. Thus, whether $\mathcal{S}_{1}$ contains a node in $\mathcal{X}$ or not, $\mathcal{S}_{1}$ is always $\gamma$-reachable.
Now we show that $\mathcal{E}$ is a minimal set. $\mathcal{G}$ has $\frac{(\gamma+1)(\gamma)}{2}$ edges to form a $(\gamma+1)$-clique. Additionally, we introduce $\gamma(\gamma-2)$ edges for connecting each of the $\gamma-2$ nodes in $\mathcal{X}^{\prime}$ with $\gamma$ nodes in $\mathcal{X}$. Summing them up, we have $\frac{3\gamma(\gamma-1)}{2}$ which equals the lower bound given in Lemma 2.
@@
2) If $\mathcal{X}\cap\mathcal{S}_{1}\neq\emptyset$, let a node $i\in\mathcal{X}\cap\mathcal{S}_{1}$. Note that after removing an edge among $\lceil\frac{\gamma-2}{2}\rceil$ pairs of nodes in $\mathcal{X}$, there exists at least one (two in case of even $\gamma$) node(s) that is connected to all other nodes in $\mathcal{G}$. In other words, $\exists v_{*}\in\mathcal{V}$ s.t. $|N_{v_{*}}|=2\gamma-1$. Now, when $1\leq|\mathcal{S}_{1}|\leq\gamma-1$, $|\mathcal{N}_{i}\setminus\mathcal{S}_{1}|\geq\gamma$, making $\mathcal{S}_{1}$ $\gamma$-reachable. In case $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$, we know there exists $v_{*}$ in either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ such that $|\mathcal{N}_{v_{*}}\setminus\mathcal{S}_{i}|\geq\gamma$, making $\mathcal{S}_{i}$ $\gamma$-reachable. Thus, whether $\mathcal{S}_{1}$ contains a node in $\mathcal{X}$ or not, either $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$ is always $\gamma$-reachable.
Now, we examine its edge set. The $\gamma$ nodes in $\mathcal{X}$ connecting to every other node adds up to a total of $\frac{\gamma(\gamma-1)+2\gamma(\gamma)}{2}=\frac{3\gamma^{2}-\gamma}{2}$ edges. Then, if we subtract $\lceil\frac{\gamma-2}{2}\rceil$ from it, we get $\frac{\gamma(3\gamma-2)+2}{2}$ for even $\gamma$ and $\frac{\gamma(3\gamma-\gamma)+1}{2}$ for odd $\gamma$, which equals the lower bound given in Lemma 5. ∎
@@
Theorem 1 provides a systematic procedure for constructing $\gamma$-MERGs, i.e., graphs that achieve maximum $r$-robustness with a minimal set of edges, thereby solving Problem 1. Examples of $\gamma$-MERGs with $n=9$ (left) and $n=10$ (right) are visualized in Figure 2. The theorem also suggests that the bounds given in Lemma 2 and 5 are tight. In fact, removing any edge from a $\gamma$-MERG immediately drops its robustness by at least one, as demonstrated later in Section 8.2.
Remark 3.
We observe that the construction of $F$-elemental graphs in [37, Prop. 1] bears similarities to our approach in Theorem 1 for odd $n$. Both methods aim to achieve maximum $r$-robustness and include a $\gamma$-clique; the key difference lies in how the remaining non-clique nodes are connected. We also observe that the total number of edges required to achieve maximum robustness for odd $n$ is the same for both approaches (see Table 1), suggesting the existence of a more general construction framework for odd $n$. While generalizing these constructions is left for future work, we note that our method additionally handles even $n$ and provides a formal proof of edge-count minimality, which is not addressed in [37].
@@ 5 Bounds on Edge Counts for $(\gamma,\gamma)$-MERGs
In Sections 3 and 4, we have examined and constructed $\gamma$-MERGs to address Problem 1. In the following two sections, we focus on $(\gamma,\gamma)$-MERGs, graphs with a given number of nodes that achieve maximum $(r,s)$-robustness with a minimal set of edges, thereby addressing Problem 2.
Similar to what we have done in the previous sections, we consider two cases: odd $n$ and even $n$. However, note that $(\gamma,\gamma)$-MERGs for odd $n$ have already been characterized in [4, Lemma 4], which is given below for completeness:
Lemma 6.
@@
Lemma 6 essentially confirms that when $n$ is odd, the graph must be a complete graph to be a $(\gamma,\gamma)$-MERG. This also implies that when $n$ is odd, for a graph to be $(\gamma,\gamma)$-robust, it must have at least $\frac{(2\gamma-1)(2\gamma-2)}{2}=(\gamma-1)(2\gamma-1)$ edges.
@@ 5.1 Graphs with Even $n$
We now introduce the lower bounds on edge counts for $(\gamma,\gamma)$-robust graphs when $n$ is even. This is formally established in Lemma 8. This lemma will be later used in Section 6 to construct $(\gamma,\gamma)$-MERGs for any $n$. However, before that, we introduce another useful lemma from [4]:
Lemma 7.
@@
Proof.
As the direct application of Equation 5, we conclude $\delta_{\min}(\mathcal{G})\geq 2(r-1)$. Then, each node has at least 2(r1)2(r-1) neighbors. Thus, since $\mathcal{G}$ is an undirected graph, $|\mathcal{E}|\geq\frac{2n(r-1)}{2}=n(r-1)$. ∎
Corollary 2 provides both the minimum degree required for each node and the total number of edges necessary for a graph to be $(r,r)$-robust. Similar to how we have used the notion of cliques as the building blocks of our analyses in Section 3, we will use Corollary 2 to analyze graphs with even $n$. By Corollary 2, we know for a graph with even $n$ to be $(\gamma,\gamma)$-MERG, it must have at least $n(\gamma-1)=2\gamma(\gamma-1)$ edges. However, by leveraging the definition of $(r,s)$-robustness (Definition 8), we derive a tighter lower bound:
Figure 3: This figure visualizes the contradiction shown in the proof for Lemma 8. The yellow solid lines represent edges, whereas the colored dotted lines represent the absences of edges. There are two disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$, in which all nodes only have $\gamma-1$ neighbors outside of the subsets, necessitating at least $\lceil\frac{\gamma}{2}\rceil$ of the colored dotted lines to become actual edges to resolve the contradiction.
Lemma 8.
@@
Proof.
By Corollary 2, we know $\delta_{\min}(\mathcal{G})\geq 2(\gamma-1)$ and $|\mathcal{E}|\geq(2\gamma)(\gamma-1)$. In the rest of the proof, we will show that some nodes need more than $2(\gamma-1)$ neighbors for $\mathcal{G}$ to satisfy definition of $(r,s)$-robustness (Definition 8). Assume to the contrary that every node $i\in\mathcal{V}$ only has $2(\gamma-1)$ neighbors and $\mathcal{G}$ is $(\gamma,\gamma)$-robust. Because (i) there are $2\gamma$ nodes and (ii) each node has $2(\gamma-1)$ neighbors, it is possible to form disjoint $\gamma$ pairs of nodes that do not have an edge together. Now, form a disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that each subset contains one node of each pair that do not have an edge. Since we have $\gamma$ pairs of nodes without an edge, $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$. By the definition of $(r,s)$-robustness, we need at least $\gamma$ nodes that have $\gamma$ or more neighbors outside of their subsets. However, we have reached a contradiction, as all nodes in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ have only $\gamma-1$ neighbors outside of their subsets. To resolve the contradiction, we need at least $\lceil\frac{\gamma}{2}\rceil$ edges, which are used to create edges between at least $\gamma$ nodes (as shown in Fig 3). Thus, combining everything, we need at least $2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$ edges. ∎
Lemma 8 provides the necessary conditions on the edge counts of $(\gamma,\gamma)$-robust graphs with even numbers of nodes. Combining Lemmas 6 and 8, we will construct $(\gamma,\gamma)$-MERGs in the next section.